OpenStudy (anonymous):

How would you integrate (ln(x^2 + 1))/(x^2) dx

OpenStudy (solomonzelman):

So I can write the question as \[\normalsize\color{blue}{ \int\limits_{ }^{ } \frac{\log(x^2+1)}{x^2}~dx} \]

OpenStudy (solomonzelman):

okay ?

OpenStudy (anonymous):

yes.

OpenStudy (solomonzelman):

you need to integrate by parts, (using § for the integral sign) § f dg = fg - §g df f= log (x²+1) df = [2x ] / [ x²+1] dg= 1/x² dx g=-1/x

OpenStudy (solomonzelman):

What I am getting here is, [ -log(x²+1) ] /[ x ] + § [ 2 ]/[x²+1] dx

OpenStudy (solomonzelman):

[ -log(x²+1) ] /[ x ] + § [ 2 ]/[x²+1] dx [ -log(x²+1) ] /[ x ] + 2§ [ 1]/[x²+1] dx [ -log(x²+1) ] /[ x ] + 2(tan^-1(x) ) + C [ -log(x²+1) ]/x + 2tan^-1(x) + C

OpenStudy (solomonzelman):

2tan^-1(x) - [ log(x²+1) ]/x + C (switching the order)

OpenStudy (solomonzelman):

It would take forever to do it in LATEX, sorry.

OpenStudy (anonymous):

thank you very much! i was thinking of using trig substitution directly and forgot about integration by parts...

OpenStudy (solomonzelman):

Anytime... also, once you mentioned trig, I'll tell you that it is the highest level course in math that I took. yw !